# skipabove=1em,skipbelow=1em

### v3

@@ 6. Effectiveness and User Acceptance of Current Ingestion Health Intervention
In this section, we examine papers that include user studies to assess the effectiveness and user acceptance of their work. To systematically understand their results, we have categorized these papers into four types: behavioral feedback/intervention, educational feedback/intervention, social feedback/intervention, and other specific feedback/intervention. Furthermore, we have divided each type into three paradigms: those that include feedback only, those that include intervention only, and those that include both feedback and intervention. In total, 106 papers are included.
6.1. Behavioral Feedback/Intervention
@@
Behavioral feedback/intervention refers to approaches that primarily focus on altering specific ingestive behaviors of users, such as eating speed, chewing counts, and drinking frequency, ultimately assisting them in achieving healthier eating habits. The technical approaches in this category primarily rely on sensor-based technologies (e.g., IMU, piezoelectric sensors, RFID) and real-time feedback mechanisms (e.g., visual, haptic, and auditory cues).
6.1.1. Behavioral Feedback Only
@@
Behavioral feedback refers to the presentation of real-time or retrospective data that directly reflects a user’s actions, enabling self-awareness and self-regulation without explicit instructions or forced behavioral changes. It objectively mirrors user behavior, allowing individuals to draw their own conclusions and adjust their habits autonomously. Systems that provide behavioral feedback only — such as continuous glucose monitoring (Liang, 2024) and smart hydration bracelets (Fortmann et al., 2014) — offer unobtrusive data visualization without explicit instructions, relying on users’ self-reflection to drive behavioral changes. Similarly, ambient logging devices (Tang et al., 2015) and smart nutrition trackers (Alshurafa et al., 2015) collect and display consumption patterns through minimal interfaces, facilitating autonomous behavior adjustment. A total of 20 papers are included in this category.
@@
Feedback effectiveness is evidenced by measurable behavioral changes. For instance, real-time glucose feedback (Liang, 2024) reduced average glucose levels in 50% of participants and prompted adjustments in meal timing, aligning with the ”action” stage of the transtheoretical model. Bite-rate feedback (Kim et al., 2016b; Scisco et al., 2011) led to a 57.7% reduction in energy intake among fast eaters, although the effects varied based on baseline consumption habits. Isomorphic feedback, such as the waist inflation from FatBelt (Pels et al., 2014) and ambient displays (Kaner et al., 2018), utilized physical or visual cues to reduce calorie intake and increase water consumption by 20–30%, respectively. Notably, the modality of feedback matters: younger users (ages 20s) preferred vibration feedback, while older adults (ages 50–70) favored light-based feedback (Kim et al., 2016b). Additionally, manual input requirements, such as the time-interval logging in the SAL logger (Tang et al., 2015), can reduce adherence, particularly among older adults.
@@
The success of behavioral feedback presumably hinges on balancing autonomy (minimal manual input) and personalization (age-appropriate feedback and aesthetic adaptability). Devices that integrate seamlessly into daily routines typically outperform those that require active engagement. Behavioral feedback demonstrates consistent patterns in user acceptance and effectiveness across multiple studies. For instance, wrist-worn devices, such as the WaterJewel bracelet (Fortmann et al., 2014) and the IDEA wristband (Lee et al., 2019), achieve a long-term usage willingness of 70–83% due to their non-intrusive design and plug-and-play functionality. In contrast, neck-worn (Fortmann et al., 2014) or head-mounted sensors face lower acceptance rates due to discomfort or social stigma (Lee et al., 2019). Additionally, aesthetic customization, such as WaterJewel’s design catering to both male and female users (Fortmann et al., 2014) and the GROW bottle’s thermo-chromic display (Kaner et al., 2018), further enhances acceptance by aligning with personal preferences.
6.1.2. Behavioral Intervention Only
@@
Behavioral interventions actively modify ingestive behaviors through direct mechanisms, such as physical constraints or real-time prompts. Physical constraint systems, like the pneumatic shape-changing spoon (Chen et al., 2022), mechanically regulate portion sizes and chewing rates by modifying the tangible interface. Digital guidance platforms, such as portion estimation interfaces (Chaudhry et al., 2016), provide prescriptive dietary recommendations via mobile apps and interactive displays. More immersive interventions, including the sensory interactive table (de Vries et al., 2020) and the smart tray system (Jung et al., 2017), combine environmental sensing with real-time behavioral prompts to actively reshape dining interactions. Unlike passive feedback that merely presents data, these interventions enforce immediate behavioral changes by altering how users interact with food. A total of 15 papers are included in this category.
@@
Behavioral intervention systems actively reshape user actions through physical, computational, or gamified means. The mandatory approach proves particularly effective for users needing structured support, as it bypasses reliance on self-motivation to drive habitual change. Interventions that physically alter user behavior demonstrate particularly strong efficacy. For instance, the Spoon’s shape-changing mechanism reduced bite size by 13.7–16.1% and total intake by 4.4–4.6% without affecting appetite (Chen et al., 2022). Additionally, intraoral closed-loop systems increased mastication force by 26% through real-time electrical stimulation (Biyani et al., 2019) and improved swallowing duration via neuromuscular triggers (Zheng et al., 2025). Structured guidance approaches have also proven effective; for example, the DIMA app enhanced portion estimation accuracy across various literacy levels (Chaudhry et al., 2016), and computer-tailored interventions achieved a twofold greater reduction in fat intake compared to improvements in physical activity through enforced rules (Vandelanotte et al., 2005). Furthermore, WaterCoaster increased drinking frequency from 2–3 to 4–6 times daily by combining personalized hydration goals with virtual rewards (Lessel et al., 2016).
@@
Regarding user acceptance of behavioral interventions, physical manipulation systems like the SSpoon (Chen et al., 2022) achieved high acceptance by subtly altering ingestive behavior without disrupting appetite. Structured guidance approaches, such as computer-tailored interventions, revealed user segmentation; simultaneous interventions were more effective for non-compliant users, while sequential approaches suited compliant individuals (Vandelanotte et al., 2005). Contextually embedded systems exhibited the widest variation in acceptance. For instance, WaterCoaster’s gamified approach engaged 70.6% users (Lessel et al., 2016), whereas the complex gamification of the Sensory Interactive Table likely led to user fatigue (de Vries et al., 2020).
6.1.3. Behavioral Feedback with Intervention
@@
Behavioral feedback with intervention combines the features of real-time data representation and mandatory guidance. It typically takes the form of comprehensive health system applications, such as the smart home-based IoT nutritional intake monitoring system (Faltaous et al., 2021), which provides unobtrusive health intervention advice through a smartphone application along with symbolic feedback mechanisms. This dual approach simultaneously offers users visualized insights while enforcing behavioral modifications through structured guidance. A total of 7 papers are included in this category.
@@
Integrated systems that combine behavioral feedback with active interventions demonstrate enhanced effectiveness and user acceptance when they balance automation with personalization. For instance, a smart home-based nutritional monitoring system achieved over 50% habit adjustment through goal-aligned mobile reminders, although privacy concerns were noted by 66.6% users (Faltaous et al., 2021). Similarly, wristband-based eating speed systems reduced bite size with a detection accuracy of 96.7%, while maintaining high acceptance through haptic or visual feedback (Kim et al., 2016a). In clinical populations, the Parkinson’s smart spoon improved stability (X/Y-axis deviation ¡15% under simulated tremors) while integrating diagnostic alerts; however, its niche design limits broader adoption (Singhal et al., 2024). Low-cost solutions, such as receipt-based nutrition tracking, proved feasible for promoting dietary awareness, particularly for monitoring calcium intake, but relied heavily on user compliance (Mankoff et al., 2002).
6.2. Educational Feedback/Intervention
@@
Educational feedback/intervention refers to feedback or interventions that have educational significance or involve educational content. This includes food and recipe recommendations, nutrition education, educational gaming, and guidance for users to adopt healthy eating habits. Ultimately, these approaches aim to enhance food literacy among users through educational approach.
6.2.1. Educational Feedback Only
@@
Educational feedback focuses exclusively on providing knowledge-based guidance to enhance learning outcomes, without direct instructional intervention prompts. It prioritizes pedagogical value over real-time behavioral modification and delivers structured information (e.g., nutritional facts, eating patterns) to cultivate cognitive awareness and self-reflection, empowering users to make informed dietary decisions. For example, a puzzle game can be utilized to enhance children’s food literacy, encouraging self-driven exploration of healthy eating patterns without explicit behavioral directives (Cai et al., 2024). Additionally, an acoustic sensor can detect ingestion and deliver real-time nutritional insights via a mobile application, thereby raising awareness of everyday dietary habits through unobtrusive monitoring (Bi et al., 2016). This type of feedback emphasizes knowledge dissemination over actual modification, leveraging playful interaction (Cai et al., 2024) or ambient data capture (Bi et al., 2016) to implicitly influence ingestive behaviors. As such, it positions itself as an educational tool for long-term nutrition self-regulation. A total of 11 papers are included in this category.
@@
Educational feedback demonstrates significant potential in improving food literacy and dietary behaviors, particularly when leveraging engaging, multi-sensory interfaces. For instance, multi-sensory puzzle games increased children’s food literacy by 59.16% through tangible interactions that fostered cognitive connections with food (Cai et al., 2024). Similarly, mixed reality (MR) food labels reduced beverage energy content by 34% and sugar content by 28% by making nutritional information more accessible and interactive (Fuchs et al., 2019). Speech-based journaling applications like FoodScrap (Luo et al., 2021) encouraged mindful eating by reducing data entry burdens, with users reporting increased self-reflection on their ingestive behaviors. In clinical populations, the SaltSwitch application (Eyles et al., 2017) successfully reduced salt purchases among patients with cardiovascular disease by providing real-time, barcode-based traffic light nutrition labels.
@@
Educational feedback excels when it reduces cognitive load (e.g., speech input in FoodScrap (Luo et al., 2021) compared to manual logging) and leverages intuitive interfaces (e.g., mixed reality labels versus traditional charts (Fuchs et al., 2019)). It achieves higher user acceptance by prioritizing intuitive interactions and minimizing user burden.
6.2.2. Educational Intervention Only
@@
Educational interventions emphasize the delivery of structured knowledge and instructional guidance through direct conversational interactions, without incorporating data-driven feedback mechanisms. Their design aligns with didactic principles, leveraging conversational interfaces to scaffold learning. This type of intervention is typically exemplified by dialogue-based chatbots, such as Foodbot (Prasetyo et al., 2020), and humanoid social robots (Robinson et al., 2021), which provide goal-oriented educational content (e.g., nutrition advice and dietary management) through scripted or AI-generated conversations. A total of 19 papers are included in this category.
@@
Social robots and embodied agents have proven particularly effective in delivering interventions, increasing intervention effectiveness by 23% compared to digital interfaces. Clinic-based humanoid robots, for instance, have shown moderate effectiveness in reducing alcohol intake while maintaining high usability (Robinson et al., 2021; Mayumi et al., 2023). Similarly, social robots outperformed chatbots in promoting healthier eating intentions among children, leveraging their physical presence to enhance persuasiveness (De Carolis et al., 2024). Interactive tableware, such as FunEat’s gamified plate (Zhao et al., 2021), improved children’s food choices through projected animations, although parents noted potential distractions during meals. Other ubiquitous computing approaches, like sensor-embedded forks, achieved 89% child engagement by replacing abstract lessons with playful interactions (Kadomura et al., 2014).
@@
Personalized educational dietary interventions demonstrate strong effectiveness when incorporating adaptive algorithms. The SmartDiet system (Hsiao and Chang, 2010) generates nutritionally balanced meal plans within seconds while maintaining high user acceptance through interactive optimization. Linear optimization frameworks have achieved less than 1.2% calorie recommendation errors by dynamically adjusting to user deviations (Arefeen et al., 2022). Intelligent nutrition services further enhance adoption by combining image recognition with flexible service selection, although ecosystem factors did not significantly impact perceived educational-targeted behavioral control (Hsiao et al., 2011). Therefore, future systems should balance guidance intensity (Liu et al., 2024) with user autonomy (Ritschel et al., 2018), while leveraging AI optimization for sustainable personalization (Arefeen et al., 2022; Hsiao and Chang, 2010).
6.2.3. Educational Feedback with Intervention
Educational feedback and intervention combine knowledge-based data insights with structured guidance to create a comprehensive learning support system. This approach delivers real-time educational feedback (e.g., nutrition analytics) while providing direct instructional interventions (e.g., personalized study plans and goal-setting prompts) to reinforce desired behaviors through a standardized framework (Kadomura et al., 2013b). Gamification is also widely applied in such scenarios (Nakamura et al., 2023). This dual mechanism ensures that learners not only understand their performance metrics but also receive actionable, curriculum-aligned strategies to bridge knowledge-practice gaps, distinguishing it from standalone feedback or intervention approaches. Related studies demonstrate a convergence of real-time monitoring and pedagogical guidance in dietary feedback and interventions. These implementations collectively operate within the ”observe-educate-act” framework through embedded sensors and algorithmic personalization, although they vary in modality (tangible vs. conversational) and intervention immediacy (real-time vs. retrospective). A total of 13 papers are included in this category.
@@ 6.2.3. Educational Feedback with Intervention
Integrated systems that combine real-time feedback with structured educational interventions achieve behavioral change through multimodal engagement, expert-guided personalization, and interactive reinforcement. Multimodal engagement proves particularly effective in pediatric contexts. For instance, systems like EducaTableware’s sound-emitting tableware (Kadomura et al., 2013b) and FoodWorks’ AR meal augmentations (Ganesh et al., 2014) increased children’s willingness to try new foods by 45-60% while reducing mealtime distractions. Meanwhile, eat2pic’s color-based food visualization significantly slowed eating speed (meal time/intervals) (Nakamura et al., 2023), although effectiveness varied with users’ artistic interests. Nutrifit’s sensor-tracked meals improved the consumption of nutrient-dense foods by 3-8% through real-time weight monitoring and app feedback (Krishna et al., 2025). Expert-guided systems also demonstrate strong clinical validity. Therefore, if such educational systems combine both embedded sensing and adaptive guidance, they will maximize user acceptance and effective outcomes.
@@
Regarding user acceptance of this combined educational approach, the OCARIoT platform reduced childhood obesity prevalence by 75.5% through IoT-collected nutrition and activity data, along with coordination from healthcare professionals, achieving over 80% user acceptance (Bastida et al., 2023). Web-based health monitoring platforms maintained 95% daily user engagement through real-time expert feedback, although 85% of nutritionists preferred weekly reviews over daily ones (Achilleos et al., 2017). Additionally, non-invasive bioelectric devices like ”Keep and Eat” demonstrated 98.45% accuracy (±1.55% error) in body composition measurements, enabling precise dietary recommendations (Balbin et al., 2020).
6.3. Social Feedback/Intervention

### v4

@@ 6. Effectiveness and User Acceptance of Current Ingestion Health Intervention
In this section, we examine papers that include user studies to assess the effectiveness and user acceptance of their work. To systematically understand their results, we have categorized these papers into four types: behavioral intervention, educational intervention, social intervention, and other specific intervention. Furthermore, we have divided each type into three paradigms: passive intervention, active intervention, and combined intervention which includes both passive and active intervention. In total, 106 papers are included.
6.1. Behavioral Intervention
@@
Behavioral intervention refers to approaches that primarily focus on altering specific ingestive behaviors of users, such as eating speed, chewing counts, and drinking frequency. The technical approaches in this category primarily rely on sensor-based technologies and real-time feedback mechanisms.
6.1.1. Passive Behavioral Intervention
@@
Passive behavioral intervention presents real-time or retrospective data reflecting a user’s actions, fostering self-awareness and autonomous behavior change without direct instruction. Systems that provide passive behavioral intervention — such as continuous glucose monitoring (Liang, 2024) and smart hydration bracelets (Fortmann et al., 2014) — offer unobtrusive data visualization without explicit instructions. Also, ambient logging devices (Tang et al., 2015) and smart nutrition trackers (Alshurafa et al., 2015) collect and display consumption patterns through minimal interfaces, facilitating autonomous behavior adjustment. A total of 20 papers are included in this category.
@@
Feedback effectiveness is evidenced by measurable behavioral changes. For instance, real-time glucose feedback (Liang, 2024) reduced average glucose levels in 50% of participants and prompted adjustments in meal timing. Bite-rate feedback (Kim et al., 2016b; Scisco et al., 2011) led to a 57.7% reduction in energy intake among fast eaters. Isomorphic feedback, such as the waist inflation from FatBelt (Pels et al., 2014) and ambient displays (Kaner et al., 2018), utilized physical or visual cues to reduce calorie intake and increase water consumption by 20–30% respectively. Notably, the modality of passive intervention matters: younger users preferred vibration feedback, while older adults favored light-based feedback (Kim et al., 2016b). Additionally, manual input requirements, such as the time-interval logging in the SAL logger (Tang et al., 2015), can reduce adherence, particularly among older adults.
@@
Passive behavioral intervention demonstrates consistent patterns in user acceptance across multiple studies. Passive behavioral intervention succeeds by balancing autonomy and personalization, with its seamless daily integration outperforming those requiring active engagement. For instance, wrist-worn devices, such as the WaterJewel bracelet (Fortmann et al., 2014) and the IDEA wristband (Lee et al., 2019), achieve a long-term usage willingness of 70–83% due to their non-intrusive design and plug-and-play functionality. In contrast, neck-worn (Fortmann et al., 2014) or head-mounted sensors face lower acceptance rates due to discomfort or social stigma (Lee et al., 2019). Additionally, aesthetic customization, such as WaterJewel’s design catering to both male and female users (Fortmann et al., 2014) and the GROW bottle’s thermo-chromic display (Kaner et al., 2018), further enhances acceptance by aligning with personal preferences.
6.1.2. Active Behavioral Intervention
@@
Active behavioral intervention actively modifies ingestive behaviors through direct mechanisms, such as physical constraints or real-time prompts. Physical constraint systems, like the pneumatic shape-changing spoon (Chen et al., 2022), mechanically regulate portion sizes and chewing rates by modifying the tangible interface. Digital guidance platforms, such as portion estimation interfaces (Chaudhry et al., 2016), provide prescriptive dietary recommendations via mobile apps and interactive displays. More immersive interventions, including the sensory interactive table (de Vries et al., 2020) and the smart tray system (Jung et al., 2017), combine environmental sensing with real-time behavioral prompts to actively reshape dining interactions. A total of 15 papers are included in this category.
@@
Active behavioral intervention systems actively reshape user actions through physical, computational, or gamified means. Interventions that physically alter user behavior demonstrate particularly strong efficacy. For instance, the Spoon’s shape-changing mechanism reduced bite size by 13.7–16.1% and total intake by 4.4–4.6% without affecting appetite (Chen et al., 2022). Additionally, intraoral closed-loop systems increased mastication force by 26% through real-time electrical stimulation (Biyani et al., 2019) and improved swallowing duration via neuromuscular triggers (Zheng et al., 2025). Mandatory approaches effectively provide structured support by circumventing reliance on self-motivation, just as structured guidance methods do; for example, the DIMA app enhanced portion estimation accuracy across various literacy levels (Chaudhry et al., 2016), and computer-tailored interventions achieved a twofold greater reduction in fat intake compared to improvements in physical activity through enforced rules (Vandelanotte et al., 2005). Furthermore, WaterCoaster increased drinking frequency from 2–3 to 4–6 times daily by combining personalized hydration goals with virtual rewards (Lessel et al., 2016).
@@
Regarding user acceptance of active behavioral interventions, physical manipulation systems like the SSpoon (Chen et al., 2022) achieved high acceptance by subtly altering ingestive behavior without disrupting appetite. Structured guidance approaches, such as computer-tailored interventions, revealed user segmentation; simultaneous interventions were more effective for non-compliant users, while sequential approaches suited compliant individuals (Vandelanotte et al., 2005). Contextually embedded systems exhibited the widest variation in acceptance. For instance, WaterCoaster’s gamified approach engaged 70.6% users (Lessel et al., 2016), whereas the complex gamification of the Sensory Interactive Table likely led to user fatigue (de Vries et al., 2020).
6.1.3. Combined Behavioral Intervention
@@
Combined behavioral intervention integrates the features of real-time data representation and mandatory guidance. It typically takes the form of comprehensive health system applications, such as the smart home-based IoT nutritional intake monitoring system (Faltaous et al., 2021), which simultaneously combines symbolic feedback with visualized insights and structured behavioral modification. A total of 7 papers are included in this category.
@@
Combined systems that integrate both passive and active interventions demonstrate enhanced effectiveness and user acceptance when they balance automation with personalization. For instance, a smart home-based nutritional monitoring system achieved over 50% habit adjustment through goal-aligned mobile reminders, although privacy concerns were noted by 66.6% users (Faltaous et al., 2021). Similarly, wristband-based eating speed systems reduced bite size with a detection accuracy of 96.7%, while maintaining high acceptance through haptic or visual feedback (Kim et al., 2016a). In clinical populations, the Parkinson’s smart spoon improved stability (X/Y-axis deviation ¡15%) while integrating diagnostic alerts; however, its niche design limits broader adoption (Singhal et al., 2024). Low-cost solutions, such as receipt-based nutrition tracking, proved feasible for promoting dietary awareness, particularly for monitoring calcium intake, but relied heavily on user compliance (Mankoff et al., 2002).
6.2. Educational Intervention
@@
Educational intervention refers to interventions that have educational significance or involve educational content for enhancing users’ food literacy. This includes food and recipe recommendations, nutrition education, educational gaming, and guidance for users to adopt healthy eating habits.
6.2.1. Passive Educational Intervention
@@
Passive educational intervention delivers pedagogical and knowledge-based guidance to enhance learning, prioritizing cognitive awareness over behavioral modification to inform dietary decisions. For example, a puzzle game can be utilized to encouraging children’s self-driven exploration of healthy eating patterns without explicit behavioral directives (Cai et al., 2024). Additionally, an acoustic sensor can detect ingestion and deliver real-time nutritional insights via a mobile application (Bi et al., 2016). This type of intervention emphasizes knowledge dissemination over actual modification, leveraging playful interaction (Cai et al., 2024) or ambient data capture (Bi et al., 2016) to implicitly influence ingestive behaviors. As such, it positions itself as an educational tool for long-term nutrition self-regulation. A total of 11 papers are included in this category.
@@
Passive educational intervention demonstrates significant potential in improving food literacy and dietary behaviors, particularly when leveraging engaging interfaces. For instance, multi-sensory puzzle games increased children’s food literacy by 59.16% through tangible interactions that fostered cognitive connections with food (Cai et al., 2024). Similarly, mixed reality (MR) food labels reduced beverage energy content by 34% and sugar content by 28% by making nutritional information more accessible and interactive (Fuchs et al., 2019). Speech-based journaling applications like FoodScrap (Luo et al., 2021) encouraged mindful eating by reducing data entry burdens, with users reporting increased self-reflection on their ingestive behaviors. In clinical populations, the SaltSwitch application (Eyles et al., 2017) successfully reduced salt purchases among patients with cardiovascular disease by providing real-time, barcode-based traffic light nutrition labels.
@@
Passive educational intervention excels when it reduces cognitive load (e.g., speech input in FoodScrap (Luo et al., 2021) compared to manual logging) and leverages intuitive interfaces (e.g., mixed reality labels versus traditional charts (Fuchs et al., 2019)). It achieves higher user acceptance by prioritizing intuitive interactions and minimizing user burden.
6.2.2. Active Educational Intervention
@@
Active educational interventions emphasize the delivery of structured knowledge and instructional guidance through direct conversational interactions. This type of intervention is typically exemplified by dialogue-based chatbots, such as Foodbot (Prasetyo et al., 2020), and humanoid social robots (Robinson et al., 2021), which provide goal-oriented educational content (e.g., nutrition advice and dietary management) through scripted or AI-generated conversations. A total of 19 papers are included in this category.
@@
Social robots and embodied agents have proven particularly effective in delivering interventions, increasing intervention effectiveness by 23% compared to digital interfaces. Clinic-based humanoid robots, for instance, have shown moderate effectiveness in reducing alcohol intake while maintaining high usability (Robinson et al., 2021; Mayumi et al., 2023). Similarly, social robots outperformed chatbots in promoting healthier eating intentions among children, leveraging their physical presence to enhance persuasiveness (De Carolis et al., 2024). Interactive tableware, such as FunEat’s gamified plate (Zhao et al., 2021), improved children’s food choices through projected animations. Other ubiquitous computing approaches, like sensor-embedded forks, achieved 89% child engagement by replacing abstract lessons with playful interactions (Kadomura et al., 2014).
@@
Personalized educational dietary interventions demonstrate strong effectiveness when incorporating adaptive algorithms. SmartDiet (Hsiao and Chang, 2010) generates nutritionally-optimized meal plans in seconds via interactive optimization, and similar approaches achieves ¡1.2% calorie errors through dynamic user adaptation (Arefeen et al., 2022). Subsequent enhancements with image recognition also improved adoption, though ecosystem factors didn’t affect educational behavior control (Hsiao et al., 2011). Therefore, future systems should balance guidance intensity (Liu et al., 2024) with user autonomy (Ritschel et al., 2018), while leveraging AI optimization for sustainable personalization (Arefeen et al., 2022; Hsiao and Chang, 2010).
6.2.3. Combined Educational Intervention
Combined educational intervention merges data-driven insights with structured guidance, delivering real-time feedback alongside instructional support through standardized frameworks (Kadomura et al., 2013b), and are often enhanced by gamification (Nakamura et al., 2023). Related studies demonstrate a convergence of real-time monitoring and pedagogical guidance in dietary interventions. This dual approach bridges knowledge-practice gaps with actionable strategies, operating via sensor-based ’observe-educate-act’ cycles across varying modalities and intervention timings. A total of 13 papers are included in this category.
@@ 6.2.3. Combined Educational Intervention
Integrated systems that add real-time feature into structured educational interventions achieve behavioral change through multimodal engagement, expert-guided personalization, and interactive reinforcement. Multimodal engagement proves particularly effective in pediatric contexts. For instance, systems like EducaTableware’s sound-emitting tableware (Kadomura et al., 2013b) and FoodWorks’ AR meal augmentations (Ganesh et al., 2014) increased children’s willingness to try new foods by 45-60% while reducing mealtime distractions. Meanwhile, eat2pic’s color-based food visualization significantly slowed eating speed (meal time/intervals) (Nakamura et al., 2023). Nutrifit’s sensor-tracked meals improved the consumption of nutrient-dense foods by 3-8% through real-time weight monitoring and app feedback (Krishna et al., 2025). Expert-guided systems also demonstrate strong clinical validity (Achilleos et al., 2017).
@@
Regarding user acceptance of this combined educational approach, the OCARIoT platform reduced childhood obesity prevalence by 75.5% through IoT-collected nutrition and activity data, along with coordination from healthcare professionals, achieving over 80% user acceptance (Bastida et al., 2023). Web-based health monitoring platforms maintained 95% daily user engagement through real-time expert feedback (Achilleos et al., 2017). Additionally, non-invasive bioelectric devices like ”Keep and Eat” demonstrated 98.45% accuracy in body composition measurements, enabling precise dietary recommendations (Balbin et al., 2020).
6.3. Social Intervention